A coin purse contains 18 coins in nickels and dimes.The coins have a total value of $1.15. Find the number of nickels and dimes.
Answer by jorel1380(3719)

o+p=18
.05o+.1p=1.15
5o+10p=115
5o+5p=90
5p=25
p=5
o=13
There are 13 nickels and 5 dimes in the purse..